### v1

# Learning Keypoints for Robotic Cloth Manipulation using Synthetic Data

Abstract
Assistive robots should be able to wash, fold or iron clothes. However, due to the variety, deformability and self-occlusions of clothes, creating general-purpose robot systems for cloth manipulation is challenging. Synthetic data is a promising direction to improve generalization, though its usability is often limited by the sim-to-real gap. To advance the use of synthetic data for cloth manipulation and to enable tasks such as robotic folding, we present a synthetic data pipeline to train keypoint detectors for almost flattened cloth items. To test its performance, we have also collected a real-world dataset.
We train detectors for both T-shirts, towels and shorts and obtain an average precision of 64.3%. Fine-tuning on real-world data improves performance to 74.2%. Additional insight is provided by discussing various failure modes of the keypoint detectors and by comparing different approaches to obtain cloth meshes and materials. We also quantify the remaining sim-to-real gap and argue that further improvements to the fidelity of cloth assets will be required to further reduce this gap. The code, dataset and trained models are available at https://github.com/tlpss/synthetic-cloth-data.
Index Terms:
Deep Learning for Visual Perception, Simulation and Animation, Data Sets for Robotic Vision
I Introduction
Clothes are everywhere in our living environments. Therefore assistive robots should be able to interact with cloth items and perform tasks such as washing, folding or ironing. However, due to their variety in shape and appearance, complex deformations and self-occlusions, cloth manipulation and perception are challenging [1, 2]. Progress has been made in recent years thanks to the continued integration of data-driven techniques [3, 4, 5] and the development of new hardware [6], but we do not yet have robots that can reliably flatten, fold or iron arbitrary pieces of cloth.
Part of the reason is that we lack the required amount of data to make these systems generalize. The use of synthetic data has proven an effective direction to overcome this data limit in various domains of robotics [7, 8, 9] and beyond [10]. Synthetic data can provide arbitrary amounts of perfectly annotated data, but the main concern is to bridge the gap between synthetic and real data, as perfectly covering the real data distribution is often infeasible. Various approaches have been explored to overcome this [7, 11, 8, 10], yet a great deal of engineering effort is usually involved in making sim-to-real work.
In this paper, we want to enable generic robot folding. To this end, we generate synthetic data to learn keypoints on various cloth items. We focus on detecting non-occluded keypoints from an RGB image of almost flattened clothes for various cloth categories. Using these keypoints, tasks such as folding flattened clothes can be achieved with scripted motions [12, 13, 5]. Several prior works use synthetic data for cloth manipulation [14, 15, 5] but to the best of our knowledge, none have tackled this specific setting. Furthermore, we are also the first to compare the performance of different procedures to obtain cloth meshes and materials.
Figure 1: In this work we learn to detect semantic keypoints on almost-flattened clothes in everyday environments. To tackle the large diversity in cloth states, cloth materials and environments, we generate synthetic data to train these keypoint detectors.
We generate synthetic data for three different cloth categories: T-shirts, towels and shorts. This process has three stages: We first procedurally create cloth meshes. In the second stage, these meshes are dropped and deformed to mimic the output of current unfolding systems, for which we use Nvidia Flex [16]. Finally, we render images of the clothes and generate the corresponding annotations. To provide real-world training data and to thoroughly evaluate the performance of the keypoint detectors, we have also gathered a dataset of almost 2,000 images containing more than 100 cloth items.
We find that keypoint detectors trained on our synthetic data have a mean average precision of 64.3 % on our test dataset, which is an improvement over the baseline trained only on real data. Fine-tuning using real data further improves performance to 74.2%. These results show that our synthetic data pipeline can be used to build state estimators for robotic cloth folding systems. Additionally, we compare several different procedures to generate cloth meshes and materials where we find that the best results are obtained by using random materials and single-layer cloth meshes, even though this decreases fidelity. Finally, we quantify the remaining sim-to-real gap and argue that this gap can only be closed by further advancements in cloth asset generation and simulation.
To summarize, our contributions are as follows:
  • We build a synthetic data pipeline and use it to train keypoint detectors that enable cloth folding.
  • We compare different procedures to obtain cloth meshes and materials, to gain more insight into synthetic data generation for cloth manipulation.
  • We provide a real-world dataset of 2,000 images and corresponding annotations of almost-flatted clothes in everyday environments.
II Related Work
II-A Robotic Cloth Manipulation
Robotic cloth manipulation has been extensively studied [1, 17, 12, 4, 3, 13]. Most works in robotic laundering focus on unfolding (also called flattening) and folding. Despite the considerable success demonstrated by unfolding pipelines [5, 6, 4], they frequently yield clothes that aren’t perfectly flattened and have yet to fully demonstrate the desired generalization across various cloth instances and environments
Given flattened clothes, tasks such as folding can be tackled using an appropriate representation and a scripted policy [5, 13]. Several methods have been explored to create such state representations for clothes, including template fitting [12, 4], edge detection [18] and semantic keypoint detection [19, 5, 14, 20]. Many of these works use depth images [14, 18, 20], as depth images are in theory invariant to cloth materials. However, they also discard useful information (such as seams) and additionally can lack the required precision [18].
In this work, we focus on learning keypoints on RGB images of almost-flattened clothes that are lying on a surface. By doing so, we aim to facilitate the creation of cloth manipulation systems that operate on the output of an unfolding system and can generalize to arbitrary clothes and environments.
II-B Synthetic data for Robotic Cloth Manipulation
Synthetic data has been used extensively in robotic cloth manipulation research to learn representations and end-to-end policies [21, 14, 15, 5, 19]. A variety of cloth simulators have been used, including Pybullet [22], Nvidia Flex[16] and Blender [23].
As always when learning from synthetic data, obtaining appropriate 3D assets is a key challenge. Many works use a limited number of pre-made cloth meshes which they manually annotate if needed [14, 5, 24]. Others generate single-layer meshes procedurally [21, 19]. The authors of [25] have focused on generating a large dataset of cloth meshes, which is used in [5, 3], though these meshes do not come with annotations of semantic locations or edges.
In this work, we further explore the use of synthetic data for keypoint detection. Our goal is to create a pipeline from which models can be trained that generalize to arbitrary clothes, deformations and environments.
III almost-ready-to-fold Clothes Dataset
TABLE I: The aRTF Clothes dataset in numbers. Both scenes and cloth items are distinct for the train and test splits to measure the generalization performance of trained models.
# scenes # cloth items # Images
Cloth Category Train Test Train Test Train Test
Towels 6 8 15 20 210 400
T-shirts 6 8 15 20 210 400
Shorts 6 8 8 9 112 180
Boxershorts 6 8 11 11 154 220
Total 6 8 49 60 686 1200
To train and evaluate the keypoint detectors, we have created a dataset of (almost) flattened clothes. We aim to enable folding clothes that have been previously unfolded (flattened) and because current unfolding pipelines do not yet produce perfectly flattened results [5, 6], we have applied small deformations to the clothes. We refer to this dataset as the aRTF Clothes dataset, for almost-ready-to-fold clothes. There have already been some efforts to collect datasets for robotic cloth manipulation [4, 26, 27], but to the best of our knowledge, this dataset is the first to provide labeled images of almost-flattened cloth items in household settings.
In the next sections, we briefly describe the data capturing and labeling procedure.
III-A Dataset Capturing
To collect the images we listed 14 household scenes containing a flat surface that can be used to fold clothes. We then collected several clothing items by asking lab members to bring them. For the towels, we included some pieces from the household cloth dataset [28].
For each image, we started by laying the cloth item on the folding surface. Based on examples from state-of-the-art unfolding pipelines [4, 5, 6], we distilled a number of failure modes and linked these to deformations that we randomly applied to the flat cloth items. These deformations include a random pinch to create wrinkles, or folding a side or corner of the cloth upwards or downwards.
We collected the images using a ZED2i RGB-D camera. For a number of scenes we used a smartphone alongside the ZED camera, to incorporate camera and lens diversity. The camera was positioned to mimic an ego-centric perspective. The distance to the folding surface is between 0.5 and 1 m.
The dataset contains a total of 1,896 images and is summarized in Table I. Collecting these images took 2 persons about 15 hours. Some example images from the dataset can be found in Fig. 1.
III-B Annotating
For each cloth category, we defined a number of semantic locations (keypoints) and annotated these on each image. We mostly use the same semantic locations as Miller et al. [29]. The semantic locations can be seen in the examples in Fig. 1.
We also labeled the segmentation mask and bounding boxes, for which we use Segment Anything (SAM) [30]. We have found SAM regularly fails with a single point, but produces satisfactory masks when provided with or 2 to 3 points on the cloth surface. In total, this labeling effort took a single person about 3 days.
IV Procedural Data Generation Pipeline
Figure 2: Examples of the generated synthetic images. Though these single-layer meshes with random materials look unrealistic, they have the best performance out of all evaluated procedures.
In this section, we describe the synthetic data pipeline we have created to generate synthetic images of clothes. In general, generating synthetic data entails two steps: obtaining the assets (meshes, materials…) and specifying how these assets should be used to create 3D scenes, from which the desired images and their corresponding annotations can then be generated. Powerful tools such as Blender[23] have enabled the latter step. The main challenge typically lies in gathering or creating the required assets, and this typically comes with a trade-off between realism and diversity. As we are dealing with deformable objects, generating the assets is even more complicated: we need not only to obtain cloth meshes but also to generate the desired distribution of configurations for them, i.e., we need to deform them. The next sections will discuss in more detail how we generate cloth meshes and materials, deform the clothes using Nvidia Flex [16], and use Blender [23] to generate images and the corresponding annotations.
IV-A Obtaining Cloth Assets
IV-A1 Cloth meshes
The first step is to obtain flat cloth meshes. We generate a series of 2D boundary vertices using a template for each cloth type. These templates are inspired on [29]. We then connect the vertices using bezier curves to create single-layer meshes. These bezier curves are used to create the neck of a T-shirt and to better mimic real cloth items by slightly curving edges. For the same reason, we also round the corners of the single-layer mesh. The parameters of the skeleton, the bezier curves and the rounding radii are sampled from manually tuned ranges. Finally, the meshes are triangulated such that they have edge lengths of at most 1cm and UV maps are generated. We keep track of the vertices that correspond to each semantic location to automatically label the keypoints later on.
IV-A2 Cloth Materials
Faithfully reproducing cloth materials requires both mimicking fabrics and producing realistic color patterns. At the same time, we also need these materials to provide sufficient diversity to enable the generalization of the models trained on the data. This combination tends to be infeasible. We have experimented with various procedures (see Section VI-B) to generate the cloth materials and have found that using random textures from PolyHaven [31] provides the best results. To increase diversity, we mix these textures with a random color.
IV-B Deforming the Meshes
To generate diverse mesh configurations from the flattened meshes, we deform them using the Nvidia Flex cloth simulator [16], which is also used in several previous works on robotic cloth manipulation [32, 3, 5]. We have also experimented with the Blender cloth simulator [23] and Pybullet [22], but found Nvidia Flex to perform better.
To deform the meshes, we randomize their orientation and drop them on a surface to create wrinkles. Afterwards, we sometimes grasp a point on the cloth and use a circular trajectory to create folds. To create both visible and invisible folds, we sometimes flip the cloth by lifting it entirely, rotating it and dropping it again. This procedure differs from the drop-pick-drop procedure used in [5, 15], as we focus on almost-flattened clothes instead of arbitrary crumpled states. The parameters of the procedure have been tuned manually to maximize performance.
To load a mesh into Nvidia Flex, we convert all vertices of the mesh to particles and add both stretch stiffness and bending stiffness constraints using the 1-ring and 2-ring neighbors of each vertex, similar to [5, 32]. Additionally, similar to previous work [3, 14, 5], we randomize various physics parameters, including bending stiffness, stretch stiffness, friction and drag to increase diversity.
TABLE II: Performance of the keypoint detector for various data sources. Training on synthetic data outperforms the baseline trained only on real data. The performance is significantly improved by training on synthetic data and then fine-tuning on real data.
mAP2,4,8 (normal-$\uparrow$) mAP2 (normal-$\uparrow$)
Cloth Type real-to-real sim-to-real (sim+real)-to-real real-to-real sim-to-real (sim+real)-to-real
T-shirt 54.4 58.2 69.1 30.5 43.8 51.9
Towel 74.4 83.2 88.6 65.9 77.7 83.5
Shorts 50.7 51.4 64.9 37.5 44.0 55.7
All 59.8 64.3 74.2 44.6 55.2 63.7
IV-C Scene Composition
Using the deformed cloth meshes and the cloth materials described before, the procedure to generate a synthetic image and its annotations is explained below.
We sample an environment texture from PolyHaven [31] to create complex scene lighting. At the same time, we use this texture to create a scene background. Thereafter, a rectangular surface is added to the scene to mimic the folding surface. We sample a material from PolyHaven [31] for this plane and randomize its base color to increase diversity and avoid a bias towards certain colors. We then put a cloth mesh on top of the table and solidify the mesh to add some thickness to the layers. Afterwards, we sample a cloth material and apply it to the mesh. To make the keypoint detectors more robust, we place a number of distractor objects on the surface. We sample the objects from the Google Scanned Objects dataset [33]. Next, we randomize the camera position in a spherical cap around the cloth center and point the camera towards the cloth. At this point we render a single image using Cycles, Blender’s Physically Based Renderer, and create a bounding box, segmentation mask and 2D keypoint annotations. For the keypoints, we use raycasting to determine if the vertex that corresponds to the keypoint is occluded. Inspired by [10], we consider a keypoint as visible if any vertex in its N-ring neighborhood is visible. This corresponds better to the human labels on the aRTF dataset.
IV-D Dataset Generation
The 3-stage procedure explained in Sections IV-B, IV-B and  IV-C is used to generate the desired amount of images for all cloth categories. Creating a mesh template is very fast. Deforming a mesh takes about 6 seconds. Rendering takes about 5 seconds for a 512x256 image on our workstation with an Nvidia RTX 3090, 32GB of RAM and an Intel i7 CPU. More details on the procedural data generation pipeline can be found in the accompanying codebase.
V Keypoint Detection
This section provides more details about the keypoint detectors. We describe the model used to detect keypoints and the average precision metrics we use to quantify the performance of the keypoint detectors. We also describe how we ordered the keypoints in image space to deal with the symmetries of cloth objects.
V-A Model and Training
We formulate keypoint detection as regression on 2D heatmaps as in [34]. With this formulation, keypoint coordinates are converted into Gaussian blobs on a 2D heatmap, and these heatmaps are then used target for pixel-wise regression. The keypoint model uses a Unet [35] inspired architecture. The encoder of the Unet is a pretrained MaxViT [36] nano model, obtained from the timm library [37].
As is common  [10, 8], we employ a set of image augmentations during training: random cropping, random Gaussian blurring and noise, and random color, contrast and brightness augmentations. Note that some of these effects could be added to the synthetic data pipeline, but we follow [10] by adding them as augmentations to increase flexibility.
We train models on synthetic and/or the train split of the aRTF dataset (see Table I). As the synthetic datasets tend to have an order of magnitude more data compared to our aRTF dataset, we use a different set of hyperparameters for each setting. These were obtained using a random hyperparameter search.
We train each model until convergence and use a subset of the aRTF train split as validation set for this purpose. The performance of each model is tested on the test split of the aRTF dataset, using the checkpoint with the best validation performance.
V-B Evaluation Metric
In this work, we want to enable folding. Hence, we only detect keypoints that are not occluded, as a robot can only interact with those parts of the cloth. This implies a varying number of detections for each keypoint category and hence we cannot directly use adistance-based metric such as the average pixel distance [14, 20]. Instead, we use the Average Precision (AP) to quantify the performance of the keypoint detectors. To distinguish between false positive and false negative keypoint predictions we use the L2 pixel distance [^1] COCO uses a non-linear transform of the L2 pixel distance: the object-similarity score (OKS). We directly use the L2 distance, since it relates more directly to real-world distances, which are typically the metric of interest for robotic manipulation..
As thresholds for the L2 distance, we use 2, 4 and 8 pixels. For the ZED 2i with an image resolution of 512x256 and given that the maximal distance between a cloth and the camera in the aRTF dataset is 1 meter (see Section I), this corresponds to real-world discrepancies between real and predicted keypoints of at most 1, 2 and 4 cm. The thresholds were chosen as follows: 1cm and 2cm are relevant for grasping as they are the half-width of regular fingertips and of custom fingertips used in [19]. 4cm is relevant for tasks such as ironing and serves as upper bound to distinguish between lack of precision and lack of semantic knowledge for the keypoint detectors. Similar in spirit to the COCO metrics [38], we report the mean average precision over all thresholds (mAP2,4,8 ) and the AP on the most strict threshold (mAP2 ).
V-C Dealing with symmetry
Clothes have (apparent) symmetries that need to be handled carefully when detecting keypoints, as the model cannot be expected to distinguish between a motion of the camera versus a motion of the cloth object. Therefore we need to handle all symmetries in the Special Euclidean group SE(3) (which describes all rigid motions and hence all possible camera movements). One possible solution is to not distinguish between symmetric keypoints by labeling them identically [19], but this results in some loss of information: we can no longer infer the connectivity of the keypoints, which is useful if you want to e.g. grasp the two corners of an edge. Other solutions include using a loss function that is invariant across all symmetries [14], or selecting an arbitrary ordering in the image [5, 24]. We have taken the latter approach in this work and have adapted the labels of the synthetic data and the aRTF dataset as follows:
For towels, we have taken the keypoint closest to the top left corner of the bounding box to be the first and its physically adjacent corner that is closest to the top right corner of the bounding box as the second keypoint. The others follow from this convention.
T-shirts or shorts have no symmetries in SE3 (which does not include reflections). However, detecting the front/back side can be very ambiguous, even on real data, and therefore we treat this as a (rotational) symmetry. For T-shirts we take as left side, the side of which the waist keypoint is closest to the bottom left corner of the T-shirt’s bounding box. For shorts, we use the top left corner of the bounding box.
VI Experiments
In this section, we train several keypoint detectors on different data sources. We evaluate the performance of all models on the test split of the aRTF dataset (see Table I) using the mean Average Precision (mAP) as metric. In Section VI-A we show that training on the synthetic data for three cloth categories results in better performance than training on the aRTF train split, and how fine-tuning on the real data improves performance significantly. The cloth mesh procedure and cloth materials described in Section 2 are compared against other alternatives in Sections VI-B and VI-C.
VI-A Main Results
Figure 3: Illustration of failure modes of the detectors. From left to right, each row shows ground truth keypoints and the output of detectors trained on real data, synthetic data and on both. The first row illustrates how the real baseline produces incomplete and inconsistent keypoints, whereas the second row shows how the detectors still struggle with folds. In the third row, the detectors are confused by an open zipper and mistake it for a leg of the shorts.
TABLE III: reality gap of the synthetic data as measured by the difference between sim-to-sim and sim-to-real performance.
mAP2,4,8 (normal-$\uparrow$) mAP2 (normal-$\uparrow$)
cloth type sim-to-real sim-to-sim sim-to-real sim-to-sim
T-shirt 58.2 81.8 43.8 69.5
Towel 83.2 84.7 77.7 78.5
Shorts 51.4 79.8 44.0 71.4
All 64.3 82.1 55.2 73.1
We have trained keypoint detectors for both T-shirts, shorts and towels. For each cloth type, we trained three models using different data sources: the corresponding train split of the aRTF dataset (which we use as a baseline), a dataset of synthetic images and a combination of both by fine-tuning on the aRTF data after training on the synthetic data. We use 10,000 synthetic images per category as we have observed that surpassing this amount does not lead to a significant improvement.
The performance of all models, as measured on the aRTF test splits, is given in Table II. Over all cloth types, the baseline model that is only trained on real data obtains an mAP2,4,8 of $59.859.8$ and an mAP2 of $44.644.6$. Training on synthetic data outperforms this baseline by 4 and 9 percentage points (pp ) respectively. Fine-tuning on the aRTF data results in an mAP2,4,8 of $74.274.2$ and an mAP2 of $63.763.7$, which is an improvement of 10 pp compared to training on synthetic data only. The performance on T-shirts and shorts is rather similar, whereas the results for towels are respectively 10 and 20 pp above average for both metrics. A few examples of the output of the models trained on a combination of the aRTF train split and synthetic data can be found in Fig. 1. Using these keypoints, folding policies can be scripted [13, 5].
The above results clearly show the added value of the synthetic data. At the same time, the results in Table II also indicate that the keypoints are still not always detected correctly, as this would result in an mAP close to 100%. To provide more insight into the failures of the detectors, we show predictions of the models trained on each data source on a number of samples from the aRTF test split in Fig. 3. Each example illustrates a common failure mode: The first row illustrates how training only on the aRTF train split results in models that confuse semantic locations and return many false positives, which is also visible in the second example. The keypoint detectors still struggle with folds, as can be seen in the second example. Despite many efforts to tailor the distribution of mesh deformations in the synthetic data (see Section IV-B), we have not managed to completely solve this problem. The final example in Fig. 3 illustrates how all models have difficulties with a particular configuration of a short in which the opened zipper is confused for a leg. Similar confusions occur for T-shirts as well.
The lack of further improvement when adding more synthetic data and the performance improvement when fine-tuning on the aRTF data already suggest that there is a reality gap for the synthetic data. To further quantify this gap, we have evaluated both the sim-to-sim performance and the sim-to-real performance of the models trained on synthetic data. These results can be found in table III. Across all categories, the gap is about 20pp , which confirms there is a substantial reality gap. Interestingly, for towels the sim and real performance are very close, suggesting that the gap is smaller for this category. This also corresponds with the results from Table II, where the performance increase associated with fine-tuning on the aRTF data was lower than for the other categories.
VI-B Comparing procedures for obtaining cloth meshes
In this experiment, we motivate the use of single-layer meshes for synthetic data generation by comparing three different procedures to generate the cloth meshes:
  1. 1.
    Inspired by [5] we have taken 5 meshes from the Cloth3D dataset [25]. We have triangulated the meshes and cleaned up their UV maps. Using Nvidia Flex, we generated a flattened version of these meshes, by dropping them from a certain height. We then manually labeled all keypoints by selecting the corresponding vertices on the mesh. Similar to [14], we have created 10 variants of each mesh by rescaling it, leading to a total of 50 different meshes. Using these meshes, we generate deformed configurations with the procedure discussed in Section IV-B.
  2. 2.
    We procedurally generate single-layer meshes using a 2D template as described in Section IV-A1. We then create deformations using the same procedure as above. As these meshes and their annotations are generated procedurally, we can generate arbitrary amounts of meshes, unlike for the cloth3d procedure.
  3. 3.
    We use the same 2D templates as above but do not deform them, which results in perfectly flat meshes, similarly to [19].
An example of each mesh procedure can be found in Fig. 4. For each procedure, we generate a dataset of 5,000 T-shirt meshes which we then use to generate 5,000 synthetic images using the pipeline described in Section IV-C. The performance on the aRTF test split of the keypoint detectors trained on these synthetic datasets can be found in table IV. Surprisingly, the more realistic meshes from the Cloth3D dataset perform worse than the less realistic but more diverse single-layer meshes. The undeformed single-layer meshes also perform surprisingly well. Based on these results, we use the single-layer mesh procedure in this paper.
Figure 4: Examples of the different cloth mesh procedures that were considered. Left to right: undeformed single-layer mesh, cloth3d mesh, single-layer mesh. Though less realistic, using the single-layer meshes results in the best performance.
TABLE IV: Comparison of different procedures to obtain cloth meshes. The single-layer procedure results in the best performance.
Cloth Mesh Procedure mAP2,4,8 (normal-$\uparrow$) mAP2 (normal-$\uparrow$)
Cloth3D subset 43.4 29.1
single-layer 54.3 38.7
single-layer undeformed 53.8 39.0
VI-C Comparing different cloth materials
In this experiment, we compare several different cloth material configurations, to motivate the use of random materials in Section 2:
  1. 1.
    For the first configuration, we simply add random materials from PolyHaven [31] to the cloth and mix the base color map with a random color to create more diversity.
  2. 2.
    In the second configuration, we apply a uniform color to the clothes as in [3, 5, 15, 19].
  3. 3.
    We also compare against a more tailored procedural material: next to uniform colors, we also add striped color maps. We also add random images to the clothes to mimic logos and prints. Furthermore, we procedurally generate a normal map that mimics fabric materials and generates additional wrinkles, similar to [19, 15].
An example of each material can be found in Fig. 5. We have generated datasets of 5,000 images for these materials, using the procedure from Section IV-C. The performance of the keypoint detectors trained on these datasets can be found in table V. We observe that uniform color materials result in the lowest performance. This is not surprising as the aRTF dataset contains many non-uniform colored cloth items. It is also consistent with findings in our previous work [19], where we also observed a decreased performance for non-uniform cloth pieces. Perhaps more surprisingly, the random materials perform better than the tailored materials even though the latter look more plausible.
Figure 5: Examples of the cloth different materials that were considered. From left to right: uniform colors, a cloth-tailored procedural material and random PolyHaven [31] textures. Though random materials are less plausible, they result in the best performance.
TABLE V: Comparison of different Cloth materials. Adding random materials to the cloth meshes leads to the highest performance on the aRTF test split.
Cloth Materials mAP2,4,8 (normal-$\uparrow$) mAP2 (normal-$\uparrow$)
PolyHaven random 54.3 38.7
cloth-tailored 49.0 36.5
uniform color 35.1 25.5
VII Discussion
We have created a synthetic data pipeline for clothes and used it to generate images and corresponding keypoint annotations of cloth items that are almost flattened. The improved performance when using this synthetic data to train keypoint detectors clearly shows its benefits. At the same time, we stress that the amount of real-world data of the baseline is limited. Scaling up real-world data collection would improve the results of the baseline. However, due to the huge diversity in cloth meshes, their configurations, cloth materials and environments, we argue that synthetic data will still be needed to obtain models that can generalize to all these aspects.
We have found that performance when training on the synthetic data increased as we traded diversity for fidelity for both cloth meshes and materials. However, we also determined that there is a reality gap that limits the sim-to-real performance. We argue that the current lack of realism in favor of diversity is a local optimum, as we do not expect more randomizations to further improve performance. Further increasing the realism of the data generation pipeline will allow to escape from the observed performance limit.
Improving performance will hence require more accurate cloth meshes that contain fine-grained elements such as seams, zippers, pockets and labels, which are all important cues to determine the semantic locations. To apply materials realistically, the meshes also need to contain accurate UV maps. Furthermore, the meshes also need to be automatically annotated to reduce engineering effort. Leveraging recent advances in generative text-to-3D models [39] could provide an exciting alternative to manually engineering all the above.
Not only do the assets need to be improved, particle-based simulators such as Nvidia Flex [16] have fundamental limitations. Next to the limited realism of the cloth-physics [3, 40], we have also found the inherent intertwining of mesh resolution with cloth thickness and deformation granularity hindering. Simulators such as C-IPC [41] provide improved realism [13], though at the cost of additional computational complexity.
Using generative models for end-to-end data augmentation as in [42] is an alternative to manual engineering of 3D assets and physics simulators, though we have found in some initial experiments that keeping the labels consistent when augmenting real data is a challenge.
Next to the sim-to-real gap, we have observed that the keypoint detectors still struggle with folds and severe deformations. This results in predictions that are between themselves not consistent and false positives on the apparent cloth edges. It is very well possible that these issues can be resolved with improved data generation, but we think that the models would benefit from more geometric reasoning capabilities. On the other hand, we are not convinced that forcing a model to predict keypoints from a static image of heavily deformed clothes lying on a surface is the best way forward: this might be an artificially hard task. Using more interactive perception as in [12] would allow the system to collect more information. Improvements to the generalization of unfolding systems could also reduce the need to handle deformed pieces in the first place.
VIII Conclusion
In this work, we have described a pipeline to generate synthetic images of clothes to advance the generalization of robotic cloth manipulation. By training a keypoint detector on images generated by this pipeline we have validated its usefulness. The resulting models and the aRTF dataset can be used to enable various tasks such as folding or ironing. Furthermore, the pipeline can be easily adapted to train models for other tasks or cloth types.
We have found that models trained on our data generation pipeline, which prioritizes diversity over fidelity, have a decreased performance when transferred to the real world. We argue that future work should focus on increasing fidelity to bridge this reality gap.
A limitation of our work is that we assume the cloth type is already known in advance, which cannot be taken for granted in real-world scenarios. We believe our data generation pipeline can be used for learning to classify clothes as well, but consider this to be out of scope.
Acknowledgments
The authors wish to thank Maxim Bonnaerens and Peter De Roovere for their valuable feedback. This project is supported by the Research Foundation Flanders (FWO) under Grant numbers 1S56022N (TL) and 1SD4421N (VDG) and by the euROBIn Project (EU grant number 101070596).